OpenStudy (anonymous):

use separation of variables to solve the initial value problem. dy/dx=(cos x)e^(y+sinx) and y=0 when x=0

myininaya (myininaya):

\[e^{-y} dy =\cos(x)e^{\sin(x)} dx\]

myininaya (myininaya):

integrate both sides

OpenStudy (anonymous):

Lol.. I don't know how to move "the x with dx" and "y with dy".. :/ can you show me that part?

myininaya (myininaya):

i multiplied dx on both sides

myininaya (myininaya):

\[e^{y+\sin(x)}=e^{y} e^{\sin(x)}\]

myininaya (myininaya):

divide both sides by e^y

myininaya (myininaya):

1/e^{y}=e^{-y}

OpenStudy (anonymous):

Oh ok. Thanks.

OpenStudy (anonymous):

How do I integrate cosxe^(sin x) dx?

myininaya (myininaya):

ley u=sin(x) => du=cos(x) dx

OpenStudy (anonymous):

? What about the e? Do I put e to the u power?

myininaya (myininaya):

\[\int\limits_{}^{}e^u du=e^u+C=e^{\sin(x)}+C\]

myininaya (myininaya):

i replace sin(x) with u and cos(x) dx with du

OpenStudy (anonymous):

Oh? But don't I integrate ? What happens with the cos...